OpenStudy (anonymous):

One day at White Cliffs, the maximum daytime temperature was 46 degrees Celsius. The overnight minimum was 57 degress Celsius colder than the maximum had been. What was White Cliffs' overnight minium temperature?

OpenStudy (butterflydreamer):

Max temperature = 46 degrees celsius Min temperature = 57 degrees celsius COLDER than the max temperature (46) Therefore: Min temperature = max - 57 = 46 - 57 = ? degrees celsius

OpenStudy (anonymous):

Thank you

OpenStudy (butterflydreamer):

nooo problem :)

OpenStudy (♪chibiterasu):

That's pretty cold . . .

OpenStudy (♪chibiterasu):

Seriously, who could go through that during the night? 0_0

OpenStudy (anonymous):

us Aussies

OpenStudy (broskishelleh):

Well, considering me living in Florida, I could give it a try

OpenStudy (butterflydreamer):

xD It never gets below 0 in Australia.. does it? O_O...

OpenStudy (♪chibiterasu):

We only get heat in Florida, the coldest we ever get is like 55-60 Farenheit in the winter. xD

OpenStudy (anonymous):

it does but you wouldn't want to give it it a try but the end of the night you will have your body heat keeping you warm

OpenStudy (♪chibiterasu):

Which is like 13-15 celsius or something.

OpenStudy (anonymous):

something like that

OpenStudy (butterflydreamer):

interesting :O I live in Australia... and i've never experienced anything below like.. 5 degrees celsius xD i think haha

OpenStudy (broskishelleh):

Huh, I didn't know there were many Openstudiers in Australia...

OpenStudy (butterflydreamer):

hurhur :3 Aussies unite!~ LOL jokes :)

OpenStudy (anonymous):

in Echuca that's were i live and Moama experience it alot, it also the way we keep warm a lot at Pukka

OpenStudy (butterflydreamer):

interesting!! I've never heard of those places :) But have a nice night!

OpenStudy (anonymous):

you too